Titration question
I pour 500ml water to 100ml 10M NaOH to dilute it between 1~2M

When I pour 500ml water, the mole would be 1.67M? (using MV=MV)

Then, I put 25ml unknown acid

Then, I put 2~3ml phenolphthalein indicator until it turns pink

How can I get volume of base used, concentration of acid from calculation, acid molarity, and percent error?

+ In this case volume of base before is 100ml and after is 600ml??
1
Expert's answer
2020-04-20T15:06:23-0400

0.1 L * 10 M = 1 mol NaOH contained already

New volume: V = 500 + 100 = 600 = 0.6 L

1 mol / 0.6 L = 1.67 M!

HA + NaOH = NaA + H2O

n (acid) = 1 mol

To get volume of base used you need to titrate the rest of it by HCl of known concentration. C1V1 = C2V2

Thus you know how much of it left, how much reacted with unknown acid, and then we could find the concentration of this acid or its molarity.

Percent error is the difference between a measured and known value, divided by the known value, multiplied by 100%